Sodium metal has a very wide liquid range, melting point at 98⁰C and boiling point at 892⁰C.                                                                   What is the liquid range in Kelvin?


1
Expert's answer
2021-03-23T08:15:14-0400
"T_1=273+98=371\ K\\T_2=273+892=1165\ K"